4 point charges Q1,Q2,Q3,Q4 are placed at 4 corners of a square 3cm on a side Find the electric potential at the center of the square.Each charge has a magnitude of 2.1 micro coulomb
1
Expert's answer
2019-05-10T11:18:56-0400
"V=V_1+V_2+V_3+V_4"

"V_1=V_2=V_3=V_4=\frac{kq}{r}"

"r=\frac{d}{\sqrt{2}}"

The electric potential at the center of the square:


"V=4\sqrt{2}\frac{kq}{d}"

"V=4\sqrt{2}\frac{(9 \cdot 10^9)(2.1 \cdot 10^{-6})}{0.03}=3.6\cdot 10^{6}\ V=3.6\ MV"